# CGRA4ML: A Framework to Implement Modern Neural Networks for Scientific Edge Computing

### v2

@@ IV-A Parameterized CGRA Engine Definition and Generation
$I$ In Channels In Cols $I_{S}{=}\lceil D_{W}/K_{H}\rceil$ $I_{T}{=}\lceil I/I_{S}\rceil$
@@
  • $R,C{=}$ Rows & columns of PE array
Figure 4: Tiling: An output matrix with a shape $[H,O]$ is generated as underlineSlices of shape $[H_{S},O_{S}]$ (a), in $HT{\times}OT$ iTerations by a CGRA. The compile-time hardware parameters: depth of weights cache, rows & columns of CGRA ($R{\times}C$) affect the runtime parameters as shown in (b).
@@
$X_{S}$ $:[N,(H_{T},H_{S}),W,(I_{T},I_{S})]$ slicing
@@
$K_{S}$ $:[K_{H},K_{W},(I_{T},I_{S}),(O_{T},O_{S})]$ slicing
@@
IV-B Dynamic Reconfiguration and Dataflow
CGRA4ML’s dataflow and tiling pattern ensures weights are maximally reused, vastly reducing off-chip data movement. The dataflow aims to optimize data reuse patterns; the weights, inputs, and output data move through the CGRA in a predictable pattern for convolutions, dense layers, and matrix multiplications, and can be modified using the CGRA4ML attributes. This dataflow is automated, and its control is decentralized across the PEs.
The Python API generates the configuration bits for a layer and passes them to the C firmware. During the setup phase, the firmware writes attributes to the AXI-Lite configuration registers. During the operation, the DMA controller reads the configuration bytes and passes them along with data through the AXI DMAs. The Weights Cache and Pixel Shifter receive these configuration bits, process them into fewer bits, and pass them along with the AXI-Streams to the CGRA. The multiplexers inside the PEs respond to these control bits and reroute the data dynamically at every clock cycle. Since the configuration moves with the data, the CGRA PEs reconfigure themselves on the fly, grouping and sharing partial data to process a layer efficiently. The runtime parameters and the precise description of our default dataflow through the CGRA are presented in Table II and Algorithm 1, respectively.
Next, we describe the two main dataflow cases: Matrix Multiplication and Convolution. The flexibility and modularity of the CGRA4ML framework enables additional use cases with minimal modification. Additional layers and activation functions are added by modifying the C firmware.
IV-B1 Matrix Multiplication
@@
Figure II shows the tiling (array partitioning), which is primarily output-stationary dataflow. When multiplying two matrices of sizes $[H,I]$ and $[I,O]$ to obtain an output matrix of size $[H,O]$, the output matrix is divided into slices of size $[H_{S},O_{S}]$ to be processed in $H_{T}{\times}O_{T}$ iterations. Each slice is processed in parallel using $R{\times}C{=}H_{S}{\times}O_{S}$ processing elements. When processing dense layers, the image batch is presented along $H$ to reuse weights across the batch.
IV-B2 Convolution
@@
The unified dataflow is designed to maximize convolution data reuse. $R$ rows of processing elements compute $R{=}H_{S}$ Height slices. The pixel shifter moves $R{+}K_{H}{-}1$ values from the AXI DMA $K_{H}$ times to exploit data locality in vertical convolution. $C$ columns of PEs re-group dynamically into $O_{S}{=}\lfloor C/K_{W}\rfloor$ groups to process one Out-Channel slice $O_{S}$, while each such group of $K_{W}$ columns process the horizontal convolution, exploiting its data locality. $H_{S}{\times}O_{S}$ output pixels are computed in parallel by the $R{\times}C$ array of processing elements. The weights cache is $C$ words wide and $I_{S}K_{H}$ rows deep. The weights stored in the weights cache are rotated and reused $O_{T}NH_{T}W$ times, maximizing the weights reuse.
IV-C Performance Analysis
@@
The number of clock cycles and the off-chip data movement required to process a layer on a CGRA with $R,C$ rows & columns of PEs are as follows:
Clock cycles $=O_{T}I_{T}(1+NH_{T}W(1+I_{S}K_{H}))$ (1)
Weight words $=O_{T}I_{T}I_{S}K_{H}C$ (2)
Input words $=O_{T}I_{T}NH_{T}WI_{S}(R+K_{H}/2)$ (3)
Output words $=NH_{\textbf{out}}W_{\textbf{out}}O.$ (4)
@@
Idle PE cols ratio ${=}[C\%K_{W}]/C{+}[O\%O_{S}]K_{W}/[CO_{T}]$ (5)
Idle PE rows ratio ${=}[H{\%}R]/H$ (6)
@@
Expression (5) is minimized by our hardware and unified dataflow design that enables the $C$ columns of CGRA to dynamically group and process Out-Channels $O$. For example, for a CGRA with $C{=}96$, the ratio of idle columns in each iteration: $(C\%K_{W})/C$ is zero for the most common layers with $K_{W}{=}\{1,3\}$, and $1\%,5\%,8\%$ for less common layers with $K_{W}{=}\{5,7,11\}$. To further minimize the overall idle ratio of columns (Exp. 5), the user can pick $C$ such that $\lfloor C/K_{W}\rfloor$ is a factor of $O$ for most layers. For ResNets, multiples of 3 and 4 ($C{=}12,24,96$), and for Transformers, multiples of 8 ($C{=}8,16,32$) are such optimal choices. Among the $R$ columns of the CGRA, Expression 6 is minimized by choosing $R$ as the common factor of the $H$ of most layers of the DNNs the user wishes to run on the generated hardware. Figure 8 demonstrates the performance efficiency and off-chip data movement when processing ResNet50 on PE arrays of different sizes.
V CGRA4ML Framework
Figure 5: Modular and extensible infrastructure of the CGRA4ML framework as described in Section V. The Python API built around Google’s qkeras, extracts the fixed-point representations of intermediate tensors, generating the binary weights and model specification. The Python frontend also generates vendor-agnostic, synthesizable SystemVerilog RTL design of the CGRA per user specifications, optimized C runtime firmware and TCL toolflows for ASIC and FPGA implementations. Our comprehensive verification suite powered by randomized, transactional, SystemVerilog testbenches interfacing with the C firmware via DPI-C, then verifies the user’s model, the generated CGRA, and runtime together as it would run on an embedded system, emulating cases like memory congestion. The CI/CD pipeline tests the entire framework on a set of models over different CGRA configurations
CGRA4ML is a modular, extensible framework with a Python frontend that processes the neural network models and generates SystemVerilog RTL of specified CGRA hardware and SoC components, TCL toolflows for FPGA and ASIC design, a production-ready C runtime firmware, and the SystemVerilog testbench suite that verifies the users’ model, generated hardware and C runtime comprehensively. This section describes the technical details of our infrastructure outlined in Fig. 5.
V-A Bundles
The Bundle is the core building block of the CGRA4ML infrastructure that acts as the gateway between the neural network built by the user and the version that runs on the hardware. In the Python front-end, Bundle is implemented as a subclass of qkeras Layer with additional functions. In the C runtime, Bundle is implemented as a set of runtime parameters and C functions that operate on them. A bundle is a set of qkeras layers that can be deterministically executed on our hardware framework, as seen in Figure 6. The layers inside a bundle can be enabled or disabled, parametrized by a set of runtime parameters, and skip connections can be optionally added. A list of such Bundles forms various architectures the scientific computing community uses, including Autoencoders, ResNets, and Transformers. Given our extensible infrastructure, the users can also extend and build their own bundles by adding to the Python front-end and C runtime.
Figure 6 demonstrates the layers in ResNet-50 grouped into a list of bundles. Some bundles only have a Conv2D layer, while others include activations, pooling, flattening, and skip connections. Our intermediate representation realizes this by optionally enabling the different layers within a bundle. The Python front-end handles the transformation, verifies the intermediate outputs by comparing with qkeras results, statically partitions them to be run on the CGRA and the CPU, and generates C firmware to execute the bundles in runtime.
Convolutional and dense layers and matrix multiplications are repetitive but computationally heavy but tensor operations that require hundreds or thousands of MAC (multiply-accumulate) to compute a single summed result. These are best suited to be accelerated in our dedicated, parameterized CGRA with $R$ rows and $C$ columns of processing elements that compute $R{\times}C(=192,768,1024)$ summed values in parallel over hundreds or thousands of clocks before storing them in on-chip memory.
Pixel-wise operations like pooling are lightweight but complex since countless edge cases need to be handled to replicate the behavior of ML frameworks like Keras. Therefore, these are executed in software on the CPU that controls the CGRA, while the engine computes the next set of summed pixels. This way, further operations can be easily added to the firmware without rewriting and verifying the hardware.
For more elaborate operations such as the attention mechanism, we provide classes that are internally built using multiple bundles and decomposed as such during the runtime. This helps the user build models such as transformers with ease.
Figure 6: Bundle: A given DNN is decomposed into a list of bundles, where a bundle is a group of layers that can be deterministically executed by the system generated by CGRA4ML. The CGRA accelerates the simple but compute-heavy operations, while the CPU executes the complex but lightweight pixel-wise operations. This flexibility allows us to add new, complex operations easily.
V-B CGRA4ML Software
The Python library is built upon qkeras for building and training quantized models. qkeras was chosen because the scientific computing community widely uses it to build and implement models on FPGAs. A bundle is implemented in Python as a subclass of a qkeras Layer, such that a list of bundles form a QModel that can be natively trained using the optimization algorithms available in qkeras [44]. qkeras supports fixed point, binary, and ternary quantization techniques, with various scaling and rounding options, such as rounding to the nearest even and stochastic rounding, to achieve numerical stability and to maximize accuracy in the smallest number of bits. Qkeras also provides quantized and hardware-friendly approximations to activation functions such as ReLu, Softmax, Sigmoid, and Tanh. During backpropagation, qkeras uses a straight-through estimator [45] to maintain stability.
After the quantization-aware training, our framework decodes the graph structure of the given neural network. The directed acyclic graph is traversed depth-first to identify dependencies between layers such output tensors and skip connections. The runtime buffer sizes for these tensors are determined based on the tensor shapes, and their lifetimes are calculated from the graph traversal to identify at which point they should be freed during the runtime. Then, each tensor in the generated runtime is statically assigned a memory address, freeing and reusing the buffers after their lifetimes. This compile-time static allocation makes the C runtime deterministic, avoiding memory leaks.
V-C CGRA4ML Firmware
CGRA4ML provides an optimized runtime written in generic C that can be compiled to ARM, RISC-V, or other targets. The Python front-end writes a config.h file that contains the model definition as a list of bundles with runtime parameters. The C runtime, compiled by the user with the model-specific config.h, handles the engine, reading inputs and writing outputs to pre-determined buffers. This allows the user to add further co-processors and Ethernet/PCIe pipelines by simply moving the data to and from these buffers.
Our firmware is built in three layers, to be modular and extensible to new architectures & embedded systems:
  • Model specification - exported from Python API
  • Runtime - independent of the architecture & model
  • Architecture specific functions
The runtime is written as a set of C functions that can be compiled and executed for any architecture (x86, ARM, RISC-V), and can be interfaced with the testbench as DPI-C functions for any simulator: Vivado XSim, Synopsys VCS, Verilator (C++). The firmware runs in parallel to the CGRA, applying pixel-wise operations to the tensor computed by the CGRA in the previous iteration. Two banks of on-chip memory are used in a ping-pong configuration. Data synchronization between the host CPU and the CGRA is achieved by acquiring an releasing the lock registers implemented in the AXI-Lite configuration register bank.
V-D CGRA4ML Verification
Verification is an often overlooked aspect of research designs, which results in significant friction when transitioning from FPGA to ASIC flows. When users implement sophisticated models and developers add new features, it is important to ensure they will work as intended on the final embedded system before investing the effort to implement them. Our verification infrastructure has been tested on Synopsys VCS, AMD-Xilinx Xsim, and Verilator (open source).
CGRA4ML offers a sophisticated verification suite that is integrated across hardware, software, and firmware. Our randomized, transactional testbench suite is built in SystemVerilog and interfaces with the IP’s AXI-Manager ports. The SystemVerilog Direct Programming Interface for C (DPI-C) is utilized in a novel manner to interface the runtime firmware functions with the SV testbench suite to achieve holistic verification.
First, a randomized tensor is passed through the user’s deep neural network model to capture intermediate and final output tensors in the software frontend. These are sliced, reshaped, and tiled to produce the expected inputs and outputs from the dataflow. During simulation, the data memory resides on the C side and is accessed from the SV side through custom functions. When the simulation begins, a function from the C runtime is called through DPI-C to load the model-specific configuration into the AXI-Lite register bank. As the simulation progresses, the AXI interfaces of the SV testbench read the data memory from the C side and feed the tensors into the RTL design. The outputs from the design are similarly stored into the memory. This process is randomized, such that the valid & ready signals of the address, data and response channels of the AXI interfaces are randomly toggled at the probability set by the user. This randomization emulates the behavior of a congested bus interface, stress-testing the system under realistic conditions and revealing bugs that would otherwise be unidentifiable. The intermediate and final results from the simulation are dumped into files, which the Python API then tests against the expected results.
Our comprehensive verification suite ensures that the model built and trained by the user works as it should on the RTL design and the firmware of the entire embedded system, before even starting the implementation process. After implementation, the same suite can be used on the gate level, time annotated netlist. This allows the developer and the user to add functionality to hardware and the runtime firmware and test it on their desktop before testing it on an FPGA, allowing agile development cycles. A set of GitHub actions running in the cloud verify the entire infrastructure using Verilator for multiple random examples after every update, enabling continuous integration and development (CI/CD), as shown in Figure 5.
V-E FPGA & ASIC Toolflows
CGRA4ML is built to generate synthesizable SystemVerilog RTL hardware to support both FPGA and ASIC implementation flows while being agnostic to vendor-specific tools and devices. Our hardware modules feature open AXI, AXI-Stream, and AXI-Lite interfaces to ensure compatibility with a wide variety of SoCs and other hardware IPs. We utilize open-source AXI and AXI-Stream cores [46, 47] in our SoC generation, modified to support ASIC flow in addition to FPGA. All these modules are wrapped into a single IP that can be easily instantiated and connected to any system with AXI ports.
V-E1 FPGA Toolflow
Our design decisions ensure that integrating the generated IP into an FPGA system is fairly straightforward. We provide modular TCL scripts to enable such integration, which automates the following process. On AMD-Xilinx FPGAs, the ZYNQ processing system is first configured to have three full AXI Subordinate ports with the maximum width (128-bit on ZCU104) for the highest throughput and one AXI-Lite Manager port for configuration. Their address mappings are updated, and the IP ports are connected to them. The TCL scripts then verify the connections, synthesize, place & route the design, exporting bitstream and generating reports.
In addition to this standard flow, the standardized AXI interfaces of CGRA4ML allow the user to integrate the generated IP into existing signal processing dataflow systems easily. Our scripts for AMD-Xilinx Vitis then create an application project, import the bitstream and C firmware, set compiler optimizations and allow the user to execute the system as baremetal.
V-E2 ASIC Toolflow
We have partnered with ARM to tape out an SoC with the IP generated by CGRA4ML integrated into NanoSoc [48], an open-source SoC platform around an ARM-M0 processor. The platform features APB ports for configuration and three AHB Subordinate ports for data movement in its expansion region. To make our design compatible with their system, we add AXI to AHB bridges around our design and a custom APB Subordinate port in place of the AXI-Lite port.
VI Results

### v3

@@ 4.4. Runtime Dynamic Reconfiguration and Dataflow
II In Channels In Cols $I_{S}{=}\lceil D_{W}/K_{H}\rceil$ $I_{T}{=}\lceil I/I_{S}\rceil$
@@
  • R,C=R,C{=} Rows & columns of PE array
Figure 8. Tiling: An output matrix with a shape $[H,O]$ is generated as underlineSlices of shape $H_{S},O_{S}]$ (a), in $HT{\times}OT$ iTerations by a CGRA. The compile-time hardware parameters: depth of weights cache, rows & columns of CGRA ($R{\times}C$) affect the runtime parameters as shown in (b).
Refer to caption Convolution through Unified Dataflow
Figure 9. Convolution through Unified Dataflow: (a) An input tensor of shape [N=1,H=6,W=6,I=3][N{=}1,H{=}6,W{=}6,I{=}3] is convolved with weights tensor of shape $K_{H}{=}5,K_{W}{=}5,I{=}3,O{=}4]$ to generate an output tensor of [N=1,H=6,W=6,O=4][N{=}1,H{=}6,W{=}6,O{=}4]. The runtime parameters (d) are determined by the dimensions of the workload, while the dimensions of the hardware determine the static parameters. In this example configuration, the R,C=4,12R,C{=}4,12 CGRA dimensions (c) map to the slice of the output tensor as shown in (a). The 12 columns of PEs dynamically group into 2 groups of 5 to process the horizontal convolutions ($K_{W}{=}5$). The two groups process two output channels ($O_{S}{=}2$) of the output tensor. The pixel shifter (b) receives 6 words of the input tensor: $R=4$ rows that correspond to the mapping in the output tensor, and $K_{H}/2{=}2$ rows below them. The last $K_{H}/2{=}2$ rows among the $R$ middle rows are stored in an on-chip SRAM to be used as the top rows when processing the next $H_{S}{=}R$ height slice. The top $K_{H}/2{=}2$ rows needed for the current $K_{H}{=5}$ vertical convolution come from this RAM. These $R{+}K_{H}-1{=}8$ words are loaded into a shift register bank and shifted 5 times. The $R$ rows of the CGRA receive the top $R$ rows of the shift register and thus obtain the vertical neighborhood needed for the convolution via shifting. After this shifting, the next input channel of the input image is loaded. After all input channels are loaded and shifted, the PEs of the CGRA (within a group) pass their accumulator values to the PE on their right to be accumulated there. The five PEs in a group (two groups are colored yellow and blue) perform the horizontal convolution this way. Since each group corresponds to an output channel, two such channels are processed in parallel.
In addition to static parametrization of the hardware architecture, cgra4ml also allows dynamic reconfiguration during runtime. Table 2 lists the runtime parameters. cgra4ml’s dataflow and tiling pattern ensures weights are maximally reused, vastly reducing off-chip data movement. The dataflow aims to optimize data reuse patterns; the weights, inputs, and output data move through the CGRA in a predictable pattern for convolutions, dense layers, and matrix multiplications, and can be modified using the cgra4ml attributes. This dataflow is automated, and its control is decentralized across the PEs.
The Python API generates the configuration bits for a layer and passes them to the C firmware. During the setup phase, the firmware writes attributes to the AXI-Lite configuration registers. During the operation, the DMA controller reads the configuration bytes and passes them along with data through the AXI DMAs. The Weights cache and Pixel Shifter receive these configuration bits, process them into fewer bits, and pass them along with the AXI-Streams as TUSER bits to the CGRA. The multiplexers inside the PEs respond to these control bits and reroute the data dynamically at every clock cycle. Since the configuration moves with the data, the CGRA PEs reconfigure themselves on the fly, grouping and sharing partial data to process a layer efficiently, without waiting to clear the pipelines. The runtime parameters and the precise description of our default dataflow through the CGRA are presented in Table 2 and Algorithm 1, respectively.
Next, we describe the two specific cases of dataflow: Matrix Multiplication and Convolution. The flexibility and modularity of the cgra4ml framework enables additional use cases with minimal modification. Additional layers and activation functions are added by modifying the C firmware.
Notation legend. $(\cdot)_{S}$ and $(\cdot)_{T}$ denote slice and iteration partitioning, respectively. We map the $R{\times}C$ PE array to output height and out-channels via $H_{S}{=}R$, $H_{T}{=}\lceil H/H_{S}\rceil$, $O_{S}{=}\lfloor C/K_{W}\rfloor$, $O_{T}{=}\lceil O/O_{S}\rceil$. Input channels are processed with parallel factor $I_{S}$ and iterations $I_{T}{=}\lceil I/I_{S}\rceil$; in our design $I_{S}$ is chosen so $K_{H}\!\cdot\!I_{S}\leq D_{W}$ (weights-SRAM depth), which gives $I_{T}{=}\lceil D_{W}/K_{H}\rceil$ used below. We reorder tensors to expose reuse: $X\!\to\!X_{T}$ (vertical locality via the pixel shifter), $K\!\to\!K_{T}$ (row-major streaming into the weights cache). Braces labeled data beats indicate DMA stream order; those labeled parallel words indicate words consumed/produced per cycle by the array.
@@
$X_{S}$ $:[N,(H_{T},H_{S}),W,(I_{T},I_{S})]\quad\text{where}\quad H_{T}{=}\lceil H/R\rceil,H_{S}{=}R,I_{T}{=}\lceil D_{W}/K_{H}\rceil,I_{S}{=}\lfloor I/I_{T}\rfloor$ slicing
@@
$K_{S}$ $:[K_{H},K_{W},(I_{T},I_{S}),(O_{T},O_{S})]\quad\text{where}\quad O_{S}{=}\lfloor C/K_{W}\rfloor,O_{T}{=}\lceil O/O_{S}\rceil$ slicing
@@
4.4.1. Matrix Multiplication
@@
Figure 2 shows the tiling (array partitioning), which is primarily output-stationary dataflow. When multiplying two matrices of sizes $[H,I]$ and $[I,O]$ to obtain an output matrix of size $[H,O]$, the output matrix is divided into slices of size $H_{S},O_{S}]$ to be processed in $H_{T}{\times}O_{T}$ iterations, such that $H_{T}{=}\lceil H/H_{S}\rceil$ and $O_{T}{=}\lceil O/O_{S}\rceil$. The rows and columns of the CGRA correspond to the size of a tile that can be processed at a time: $R{\times}C{=}H_{S}{\times}O_{S}$. When processing dense layers, the image batch is presented along $H$ to reuse weights across the batch.
4.4.2. Convolution
The unified dataflow is designed to maximize convolution data reuse. Fig. 9 describes a convolution workload through the unified dataflow with an example workload. $R$ rows of processing elements compute $R{=}H_{S}$ Height slices. Algorithm 1 describes the preparation of the input, weights and output tensors to be processed by the CGRA. The reuse-aware bandwidth model in Eqs. 1-4 of Section 6.1 quantitatively matches the schedule above and guides $R,C,I_{S}$ choices explaining the utilization/data-movement trends in Fig.12.
An input tensor $X$ of shape $[N,H,W,I]$ is sliced according to the static parameters and reshaped into a tensor $X_{T}$ of shape $I_{T},N,H_{T},W,I_{S},R{+}K_{H}{-}1]$. Since the output side has much lower bandwidth, slicing and reshaping the input for the next layer happens at the output side of the current layer, and the words are written to the respective locations in memory to form the tensor $X_{T}$. When processing the next layer, the DMAs read $X_{T}$ in row-major order (as written). The pixel shifter takes $R{+}K_{H}{-}1$ values from the AXI DMA and shifts them, reducing the required input bandwidth by $K_{H}\times$ by exploiting data locality.
The weights cache has two buffers that works in ping-pong configuration, each $C$ words wide and $D_{W}$ rows deep. The weights tensor $K$ of shape $K_{H},K_{W},I,O]$ is sliced and reshaped into a tensor $K_{T}$ of shape $I_{T},O_{T},I_{S},K_{H},O_{S},K_{W}]$. This slicing and reshaping is done offline, by the Python frontend, which then generates the weights binary that is loaded with the firmware, and read by the DMA in row-major order (as is). The weights stored in the weights cache are rotated and reused $O_{T}NH_{T}W$ times, maximizing the weights reuse and reducing the required bandwidth proportionally.
@@ 4.4.2. Convolution
$C$ columns of PEs re-group dynamically into $O_{S}{=}\lfloor C/K_{W}\rfloor$ groups to process one Out-Channel slice $O_{S}$, while each such group of $K_{W}$ columns process the horizontal convolution, exploiting its data locality, and reducing input bandwidth by the same amount. $H_{S}{\times}O_{S}$ output pixels are computed in parallel by the $R{\times}C$ array of processing elements.
Taken together, these components form a robust and extensible foundation for deploying quantized neural networks in embedded scientific systems. By abstracting hardware complexity through modular design and automation, cgra4ml allows scientists and developers to focus on model innovation without being encumbered by low-level engineering overhead. The ability to generate synthesizable RTL, runtime firmware, and verified execution flows ensures consistency across simulation, emulation, and deployment platforms. Moreover, the framework’s flexibility in supporting different processor architectures, SoC shells, and toolchains makes it well-suited to the rapidly evolving demands of scientific edge computing. The seamless integration of verification infrastructure, programmable CGRA engines, and composable bundles underscores the unique value of the framework as a research and deployment tool for ML-powered embedded systems.
5. cgra4ml: Verification and Deployment
Robust verification and deployability are crucial for closing the loop between neural network design and reliable hardware execution, particularly in scientific edge applications, where correctness, reproducibility, and system integration are paramount. While many accelerator frameworks focus on performance or programmability, cgra4ml places equal emphasis on comprehensive verification and streamlined hardware deployment. This section describes the verification methodology that ensures functional consistency across model, firmware, and RTL, followed by the deployment toolflows for FPGA and ASIC targets. Our infrastructure supports agile development and testing through simulation, CI/CD automation, and real-world deployment on programmable SoCs and standard ASIC flows, giving researchers and developers confidence that their designs will perform reliably across the entire development stack.
5.1. Comprehensive Verification Approach
Refer to caption Verification Infrastructure developed for cgra4ml
Figure 10. Verification Infrastructure developed for cgra4ml. The subsystems of CGRA, cache, DMAs and DMA controller are described in SystemVerilog, and during verification are connected to the production-ready C/C++ firmware through AXI-specific VIPs and custom DPI-C wrappers. This way, the same firmware that runs on the FPGA also runs in simulation, verifying the entire system holistically.
Verification is an often overlooked aspect of research designs, which results in significant friction when transitioning from FPGA to ASIC flows. When users implement sophisticated models and developers add new features, it is important to ensure they will work as intended on the final embedded system before investing the effort to implement them. Our verification infrastructure has been tested on Cadence Xcelium, AMD-Xilinx Vivado Xsim, and Verilator (open source).
cgra4ml provides a verification suite that is integrated across hardware, software, and firmware, targeting 100% bit parity with a CPU/GPU golden run at the layer, bundle, and model outputs, using the exact numeric format used by the hardware. Our randomized, transactional testbench suite is built in SystemVerilog and interfaces with the IP’s AXI-Manager ports. The valid and ready probabilities define the randomness in toggling the s_valid and m_ready handshake signals of the AXI interfaces. Setting the probabilities to 1 gives the fastest simulation, where data is transferred every cycle. This is helpful as a smoke test when a user intends to modify the hardware and debug their modification. Setting it to lower values, e.g., 0.1 and 0.01, stress tests the hardware using randomized transactions, exposing hardware bugs that smoke tests cannot reveal. The SystemVerilog Direct Programming Interface for C (DPI-C) is utilized in a novel manner to interface runtime firmware functions with the SV testbench suite to achieve holistic verification.
From the user’s perspective, this verification process is streamlined as follows. First, a randomized tensor is passed through the user’s deep neural network model to capture intermediate and final output tensors in the software frontend. These are sliced, reshaped, and tiled to produce the expected inputs and outputs from the dataflow. During simulation, the data memory resides on the C side and is accessed from the SV side through custom functions. When the simulation begins, a function from the C runtime is called through DPI-C to load the model-specific configuration into the AXI-Lite register bank. As the simulation progresses, the AXI interfaces of the SV testbench read the data memory from the C side and feed the tensors into the RTL design. The outputs from the design are similarly stored into the memory. This process is randomized, such that the valid & ready signals of the address, data and response channels of the AXI interfaces are randomly toggled at the probability set by the user. This randomization emulates the behavior of a congested bus interface, stress-testing the system under realistic conditions and revealing bugs that would otherwise be unidentifiable. The intermediate and final results from the simulation are dumped into files, which the Python API then tests against the expected results.
Our comprehensive verification suite ensures that the model built and trained by the user works as it should on the RTL design and the firmware of the entire embedded system, before even starting the implementation process. All simulations are checked against a CPU/GPU golden reference with identical datatype to the RTL (same data type, per-tensor scale/zero-point, padding/stride/dilation, and the RTL’s rounding and saturation rules). The same checks can be executed on the FPGA build by streaming the identical tensors through the AXI DMA and comparing the captured outputs to the golden results, enabling end-to-end parity checks beyond simulation. After implementation, the same suite can be used on the gate level, time annotated netlist. This allows the developer and the user to add functionality to hardware and the runtime firmware and test it on their desktop before testing it on an FPGA, allowing agile development cycles. A set of GitHub actions running in the cloud verify the entire infrastructure using Verilator for multiple random examples after every update, enabling continuous integration and development (CI/CD), as shown in Figure 6.
5.2. FPGA and ASIC Deployment
cgra4ml is built to generate vendor-agnostic synthesizable SystemVerilog RTL hardware components and necessary scripts to support both FPGA and ASIC implementation flows with various tools and targets.
5.3. FPGA Toolflow
Our design decisions ensure that integrating the generated IP into an FPGA system is fairly straightforward. Modern FPGAs are often Programmable System-on-Chips, where on-chip hardened processors and the FPGA fabric are interconnected to deploy heterogeneous systems. We provide modular TCL scripts to enable such integration, which automates the following process. On AMD-Xilinx FPGAs, the ZYNQ processing system is first configured to have three full AXI Subordinate ports with the maximum width (128-bit on ZCU104) for the highest throughput and one AXI-Lite Manager port for configuration. Their address mappings are updated, and the IP ports are connected to them. The TCL scripts then verify the connections, synthesize, place & route the design, exporting bitstream and generating reports.
In addition to this standard flow, the standardized AXI interfaces of cgra4ml allow the user to easily integrate the generated IP into existing signal processing dataflow systems. Our scripts for AMD-Xilinx Vitis then create an application project, import the bitstream and C firmware, set compiler optimizations, and allow the user to execute the system as baremetal.
5.4. ASIC Toolflow
We have partnered with ARM to tape out an SoC with the IP generated by cgra4ml integrated into NanoSoC (David Flynn, 2023), an open-source SoC platform around an ARM-M0 processor. The platform features APB ports for configuration and three AHB Subordinate ports for data movement in its expansion region. To make our design compatible with their system, we add AXI to AHB bridges around our design and a custom APB Subordinate port in place of the AXI-Lite port.
Together, the verification and deployment flows provided by cgra4ml form a production-grade pipeline that supports model prototyping, debugging, and deployment in a unified environment. By enabling cycle-accurate testing with DPI-C integration, runtime configuration emulation, and full-stack simulation, the framework significantly reduces the time and risk associated with edge ML system development. Furthermore, the modular FPGA and ASIC toolchains allow seamless migration from early experimentation to large-scale integration or even silicon realization. Whether targeting a PYNQ-based research prototype or an industrial-grade SoC, cgra4ml ensures that hardware, software, and machine learning models operate in concert, providing a reliable foundation for next-generation scientific computing at the edge.
5.5. Ibex SoC Integration
To demonstrate end-to-end deployability on an open platform, we integrated cgra4ml with the lowRISC Ibex RISC-V SoC, which we build and simulate via the FuseSoC flow. We add an AXI crossbar to unify the engine’s three AXI managers into a single port, simplifying memory mapping and arbitration. We also add lightweight bridges that translate between the Ibex load/store interface and AXI, allowing the CGRA engine to present a device port for configuration and one host port to the system. The subsystem is then packaged as a reusable FuseSoC IP core. The hardware/firmware interface is platform agnostic. On the software side, we compile our runtime and model firmware with a standard RISC-V toolchain, load the ELF together with quantized weights into on-chip SRAM, and execute the workload on the full SoC. This integration validates cgra4ml ’s hardware/firmware stack in a fully open, reproducible RISC-V environment and provides a template for broader open-source SoC deployments.
6. Experimental Evaluation
6.1. Performance Analysis and Practical Guidelines
@@
The performance of the CGRA for a given bundle can be analyzed based on the runtime parameters ($N,H,H_{S},H_{T},W,I,I_{S},I_{T},O,O_{S},O_{T},K_{W},K_{H}$) described in Table 2. The number of clock cycles and the off-chip data movement required to process a layer on a CGRA with $R,C$ rows & columns of PEs are as follows:
(1) Clock cycles $=O_{T}I_{T}(1+NH_{T}W(1+I_{S}K_{H}))$
(2) Weight words $=O_{T}I_{T}I_{S}K_{H}C$
(3) Input words $=O_{T}I_{T}NH_{T}WI_{S}(R+K_{H}/2)$
(4) Output words $=NH_{\textbf{out}}W_{\textbf{out}}O.$
@@
(5) Idle PE cols ratio ${=}[C\%K_{W}]/C{+}[O\%O_{S}]K_{W}/[CO_{T}]$
(6) Idle PE rows ratio ${=}[H{\%}R]/H$
@@
Expression (5) is minimized by our hardware and unified dataflow design (Algorithm 1) that enables the $C$ columns of CGRA to dynamically group and process Out-Channels $O$. For example, for a CGRA with C=96C{=}96, the ratio of idle columns in each iteration: $(C\%K_{W})/C$ is zero for the most common layers with $K_{W}{=}\{1,3\}$, and $1\%,5\%,8\%$ for less common layers with $K_{W}{=}\{5,7,11\}$. To further minimize the overall idle ratio of columns (Exp. 5), the user can pick $C$ such that $\lfloor C/K_{W}\rfloor$ is a factor of $O$ for most layers. For ResNets, multiples of 3 and 4 (C=12,24,96C{=}12,24,96), and for matrix multiplication based workloads, multiples of 8 (C=8,16,32C{=}8,16,32) are such optimal choices. Among the $R$ columns of the CGRA, Expression 6 is minimized by choosing $R$ as the common factor of the $H$ of most layers of the DNNs the user wishes to run on the generated hardware. Figure 12 demonstrates the performance efficiency and off-chip data movement when processing ResNet50 on PE arrays of different sizes.
[…]
Figure 11. Scaling of hardware resource and power consumption on a Xilinx ZCU104 FPGA, for different bitwidths and PE columns, with PE rows = 8
Refer to caption Performance analysis of ResNet-50
Figure 12. Performance analysis of ResNet-50 on different CGRA array sizes. Performance efficiency measures how well each layer is dynamically mapped over statically configured hardware. While larger arrays reduce off-chip data movement due to high data reuse, they have lower utilization for smaller layers. The unified dataflow of the array (Sec. 4.4) maximizes the utilization to near 100% and minimizes data movement for most layers.
Model ResNet-50 PointNet (Qi et al., 2017) Autoencoder (Team, 2025c)
Bits 4 4 8
Operations 5.36 billion 73.7 million 2.09 million
PEs (7,96) (32,32) (16, 64)
Frequency (MHz) 250 250 250
$F$s 101706 69277 154247
LUTs 82200 100076 140357
BRAMs 6 4.5 4.5
Static Power (W) 0.700 0.700 0.702
Dynamic Power (W) 3.847 3.840 4.196
Total Power (W) 4.547 4.540 4.898
GOPs/W 37.3 56.8 1
Table 3. Implementation of ResNet-50 and Pointnet on ZCU104 FPGA